A car travels 540 km in 4.5 hours. How far will it go in 8 hours at the same average speed?


1
Expert's answer
2021-05-27T08:31:53-0400

"s=vt,\implies s\sim t,"

"s_2=s_1\frac{t_2}{t_1},"

"s_2=540\cdot\frac{8}{4.5}=960~km."